A shopper pushes a 7.8- kg shopping cart up a 13 degree incline. Find the magnitude of the horizontal force, F⃗, needed to give the cart an acceleration of 1.25 m/s2.
Answer #1

Here,

m = 7.8 Kg

Using second law of motion

m * a = net force

7.8 * a = F * cos(13 degree) - m * g * sin(theta)

7.8 * 1.25 = F * cos(13 degree) - 7.8 * 9.8 * sin(13 degree)

solving for F

F = 27.7 N

the magnitude of horizontal force is 27.7 N